Assuming that banks choose to maintain a liquidity ratio of 20 per cent and assuming that new cash deposits of $100m are made in the banking system described in the table below:

$m

$m

Banks receive

100

Hold

Lend

20

80

Second round deposits rise by

Hold

Lend

Third round deposits rise by

Hold

Lend

Fourth round deposits rise by

   

Hold

Lend

   

   

Fifth round deposits rise by

     

Hold

Lend

     

  

Total deposits after five rounds

    (a)    Complete the above displayed table which shows how credit is created. (5 marks).

    (b)   How much credit will have been created after five rounds? (1 mark)…………………..

    (c)    To what level will total deposits eventually increase? (1 mark)………………………..

   

    (d)   Define the bank multiplier. (1 mark)…………………………………………………..

   

    (e)    What is the bank multiplier in this case? (1 mark)...........................................................

   

    (f)    How is it related to the liquidity ratio? (1 mark)..............................................................
Answer #1

A.

$m
$m
Banks receive 100 Hold 20
Lend 80
Second round deposits rise by 80 Hold 16
Lend 64
Third round deposits rise by 64 Hold 12.80
Lend 51.20
Fourth round deposits rise by 51.2 Hold 10.24
Lend 40.96
Fifth round deposits rise by 40.96 Hold 8.19
Lend 32.77
Total deposits after five rounds 336.16

B

Total credit created after five rounds = 80+64+51.2+40.96+32.77

Total credit created after five rounds = $268.93 M

C.

Total deposit increase = 100*(1/20%)

Total deposit increase = $500 M

D.

It refers to the number of times the money is created when $1 is deposited in the bank.

E.

Bank multiplier = 1/required reserve ratio = 1/20%

Bank multiplier = 5

F.

A higher multiplier, increase more money creation and it also increases the current assets. As a result, liquidity ratio increases.